An executive invests $23,000, some at 6% and some at 4% annual interest. If he receives an annual return of $1300, how much is invested at each rate?
at 6% interest $
at 4% interest $

Answer by macston(5194)

.
F=amount at 4%; S=amount at 6%
.
F+S=$23000
F=$23000-S
.
0.04F+0.06S=$1300
0.04($23000-S)+0.06S=$1300
$920-0.04S+0.06S=$1300
0.02S=$380
S=$19000
.
F=$23000-S=$23000-$19000=$4000
.
ANSWER: He invested $4000 at 4% and $19000 at 6%.
.
CHECK:
0.04F+0.06S=$1300
0.04($4000)+0.06($19000)=$1300
$160+$1140=$1300
$1300=$1300